OpenStudy (anonymous):

how do i multiply integers and decimals like this one. -7 multiplied by 8

OpenStudy (anonymous):

multiply the two numbers and the if the sign was (-)*(+)=(-) (+)*(-)=(-) (-)*(-)=(+) (+)*(+)=(+)